OpenStudy (anonymous):

Find the volume of the region enclosed by z=1-y^2 and z=y^2-1 for 0<=x<=54 I got : int (8-8y^2)dy

OpenStudy (anonymous):

I got : int (8-8y^2)dy eval from 0 to 54 but it's wrong

OpenStudy (anonymous):

Give me a second :P

OpenStudy (anonymous):

k thanks malevolence, I wish I could keep you on retainer, haha

OpenStudy (anonymous):

\[\int_0^{54}\int_{-1}^{1}(1-y^2) - (y^2 - 1) dydx \]

OpenStudy (anonymous):

I think that's how you start it anyway.

OpenStudy (anonymous):

haha wow, I was way off

OpenStudy (anonymous):

\[= 108\int_{-1}^1 1-y^2dy\]

OpenStudy (anonymous):

We find the bounds on y by finding the y values for where the two curves intersect: \[1-y^2 = y^2-1 \implies 2y^2 = 2 \implies y = \pm 1\]

OpenStudy (anonymous):

We are given the bounds on x.

OpenStudy (anonymous):

yup 144

OpenStudy (anonymous):

was right

OpenStudy (anonymous):

I used the wrong bounds for y and I think I did the first intergal wrong, haha. Thanks for the help polpak!

OpenStudy (anonymous):

Sure thing